# SlumpGuard: An AI-Powered Real-Time System for Automated Concrete Slump Prediction via Video Analysis

### v1

@@ Abstract
Concrete workability is essential for construction quality, with the slump test being the most common on-site method for its assessment. However, traditional slump testing is manual, time-consuming, and prone to inconsistency, limiting its applicability for real-time monitoring. To address these challenges, we propose SlumpGuard, an AI-powered, video-based system that automatically analyzes concrete flow from the truck chute to assess workability in real time. Our system enables full-batch inspection without manual intervention, improving both the accuracy and efficiency of quality control. We present the system design, a the construction of a dedicated dataset, and empirical results from real-world deployment, demonstrating the effectiveness of SlumpGuard as a practical solution for modern concrete quality assurance.
@@
journal: Journal of Building Engineering
@@
[label1]organization=Yonsei University,addressline=Yonsei University 50 Yonsei-ro, Seodaemun-gu, city=Seoul, postcode=03772, country=Republic of Korea
@@ 1 Introduction
$\dagger$\$$\dagger$\$footnotetext: Equal contributionfootnotetext: Email address: winston1214@yonsei.ac.kr (Youngmin Kim)
Concrete is the most widely used construction material worldwide, forming the structure of buildings, infrastructure, and civil engineering projects. Among the various properties of fresh concrete, workability is a key parameter, as it directly influences the ease of mixing, transporting, placing, and compacting concrete on site [1, 2]. The slump test, introduced in the early 20th century, remains the most widely adopted method for assessing concrete workability in the field due to its simplicity and low cost.
Despite its ubiquity, the traditional slump test suffers from several limitations. Its accuracy is highly dependent on operator skill, environmental conditions, and proper sampling, leading to potentially inconsistent or non-representative results [1]. Furthermore, as a manual procedure, it is inherently time-consuming and unsuitable for continuous or real-time quality monitoring, especially in large-scale construction scenarios that demand rapid decision-making [2, 3].
To address the limitations of trainditional slump testing, we introduce SlumpGuard, a video-based system for real-time monitoring of concrete workability on construction sites. By analyzing the flow of ready-mix concrete from the truck chute, SlumpGuard enables automated, full-batch inspection of every delivery without manual intervention. This approach not only enhances the consistency and reliability of slump measurements, but also streamlines quality assurance procedures across the entire construction workflow. Ultimately, SlumpGuard contributes to safer, more efficient, and higher-quality construction outcomes by providing continuous and objective assessment of fresh concrete properties.
In this study, we present an AI-powered automated system for concrete quality control, along with the construction of a dedicated dataset for model training. We further demonstrate the effectiveness of our approach through empirical evaluations and real-time on-site implementation.
2 Preliminary
@@ 2.1 Optical Flow
Optical flow is a computer vision technique that tracks the movement of pixels in an image over time to analyze the motion of objects or the camera. It is primarily used to estimate the motion of objects or scenes in video or continuous image sequences, representing the speed at which each pixel in the image changes, which is expressed as a vector field corresponding to the movement of the camera or objects.
There are two main methods for calculating optical flow: dense optical flow, which computes the flow for every pixel, and sparse optical flow, which computes the flow only for selected feature points. Sparse optical flow, while computationally more efficient, may fail to capture the full motion of the video. In contrast, dense optical flow [4] provides a more detailed representation of motion across the entire image, but it is slower. In this paper, considering the need for real-time performance and the goal of tracking concrete movement, we employ the Lucas-Kanade [5] method, which calculates a sparse optical flow.
2.1.1 Lucas-Kanade Method
@@
The Lucas-Kanade method calculates optical flow within a small window based on a local linear approximation. This method assumes that the intensity of the same point remains constant over time.
$I(x,y,t)=I(x+\Delta x,y+\Delta y,t+\Delta t).$ (1)
Based on this assumption, optical flow is calculated by performing an approximation using derivatives. This can be approximated in the following form using a Taylor series expansion.
$\frac{\partial I}{\partial x}u+\frac{\partial I}{\partial y}v+\frac{\partial I}{\partial t}=0.$ (2)
Here, $u$ and $v$ represent the optical flow in the $x$ and $y$ directions, respectively. Based on this equation, a single equation is provided for each pixel, and by using a small window, multiple equations are solved to estimate the optical flow.
2.2 You Only Look Once (YOLO)
You Only Look Once (YOLO) [6] is a real-time object detection framework that formulates object detection as a single-stage regression problem. Unlike traditional two-stage detectors such as R-CNN [7] and Fast or Faster R-CNN [8, 9], which separate region proposal and classification, YOLO directly predicts bounding boxes and class probabilities from an input image in a single forward pass through a deep neural network. This approach enables high-speed inference while maintaining competitive accuracy.
YOLO employs a convolutional neural network (CNN) to process the entire image at once. The network divides the image into the $S\times S$ grid, where each grid cell predicts a fixed number of bounding boxes along with confidence scores and class probabilities. Specifically, each cell predicts $B$ bounding boxes, each represented by coordinates $(x,y,w,h)$, an objectness score, and $C$ class probabilities. This design allows YOLO to capture spatial dependencies and global context within an image efficiently.
The loss function of YOLO consists of three main components: localization loss, confidence loss, and classification loss. The localization loss penalizes errors in the predicted bounding box coordinates using mean squared error (MSE). The confidence loss ensures that the predicted objectness scores align with the presence or absence of an object in the bounding boxes. Finally, the classification loss minimizes the discrepancy between predicted and ground-truth class probabilities using cross-entropy loss.
Building upon these model innovations, recent versions of YOLO [10, 11] have further evolved, introducing enhanced capabilities such as oriented object detection [12]. This extension is particularly useful for applications like aerial imagery and scene text detection, where objects may appear at arbitrary angles. By predicting rotated bounding boxes, these models achieve more precise localization, improving YOLO’s versatility in handling objects that are not aligned with the image axes.
2.3 ResNet 3D
ResNet 3D [13] is a model architecture designed to process video by simultaneously capturing spatial and temporal information. It has been primarily explored in two architectural variants: the Mixed Convolution structure and the (2+1)D convolutional block.
The Mixed Convolution design applies 3D convolutions only in the early layers of the network, followed by 2D convolutions in the higher layers. This design is based on the idea that motion modeling is best handled at lower levels of the network, while spatial reasoning over motion features can be efficiently performed by 2D convolutions at later stages. MC ResNets have been shown to achieve 3–4% higher clip-level accuracy compared to 2D ResNets of similar capacity, and they match the performance of full 3D ResNets despite having significantly fewer parameters.
The (2+1)D convolutional block factorizes a 3D convolution into a 2D spatial convolution followed by a 1D temporal convolution. This separation introduces an additional nonlinearity between the two operations, effectively increasing the representational capacity of the model. Moreover, this decomposition facilitates optimization, often resulting in lower training and testing losses compared to full 3D convolutions.
@@ 3.2 Preprocessing and Annotation
Since we employed a stereo camera setup, the collected videos consisted of two separate streams. Therefore, as part of preprocessing, the original stereo data was divided into two individual video sequences, effectively doubling the amount of data for analysis. Considering that concrete discharges directly from the ready-mixed concrete chute, our preprocessing specifically targeted the chute region. Bounding box annotations were performed by categorizing into two classes, as shown in Figure 2, an “Unrotated Chute” representing the entire chute region with axis-aligned boxes, and a ”Chute” represented by rotated bounding boxes.
@@ 4 Automation Strategy
    Assessing whether the chute’s discharged concrete slump falls within a specified range.
@@ 4.1 Chute Detection
To automatically identify when concrete is falling from the chute, we first detect the chute region in the video using an object detection model. We employ YOLOv8 [14], a state-of-the-art object detection model, known for its fast inference performance and high accuracy, making it suitable for real-time object detection. We initialize the model with weights pretrained on the DOTA dataset [15], which is designed for detecting oriented objects in aerial imagery and is well-suited for our task of detecting rotated chutes. Using this model, we simultaneously predict both axis-aligned and rotated bounding boxes for the chutes, enabling us to detect the chute area in pouring videos. The detected chute is represented by a bounding box with parameters $(x,y,w,h,\theta)$, indicating the center coordinates, width, height, and rotation angle of the box.
@@ 4.2.1 Definition of the Bottom Edge Direction
$\mathbf{p_{1}}=\begin{bmatrix}x\\ y\end{bmatrix}+\frac{w}{2}\begin{bmatrix}\cos(\hat{\theta})\\ \sin(\hat{\theta})\end{bmatrix},\quad\mathbf{p_{2}}=\begin{bmatrix}x\\ y\end{bmatrix}-\frac{w}{2}\begin{bmatrix}\cos(\hat{\theta})\\ \sin(\hat{\theta})\end{bmatrix}$ (4)
4.2.2 Determining the Location and Timing.
To make the temporal change explicit, we define the center point of the tracked chute in each frame as $(x_{t},y_{t})$ at time $t$, and as $(x_{t-1},y_{t-1})$ at the previous frame $t-1$. The optical flow vector $(u,v)$ represents the displacement from frame $t-1$ to $t$, so that $(x_{t},y_{t})=(x_{t-1}+u,y_{t-1}+v)$. The endpoints of the bottom edge of the rotated bounding box, $\mathbf{p_{1}}=(x_{1},y_{1})$ and $\mathbf{p_{2}}=(x_{2},y_{2})$, are computed.
To determine whether the tracked center point crosses the bottom edge between frames $t-1$ and $t$, we explicitly check the positions at both time steps relative to the edge. First, the line equation for the bottom edge is:
@@ 4.3.1 Data Preprocessing
We crop the unrotated bounding box corresponding to $V$ from the full image to isolate the chute area. From the frame at time $t$, we extract a sequence of $N$ consecutive frames from the cropped region to construct a temporally continuous input.
@@ 4.3.2 Data Augmentation
Together, these augmentations help the model handle diverse visual appearances–such as changes in brightness, shadows, occlusion, and orientation–that are commonly encountered in outdoor construction environments. They also promote better regularization by expanding the training distribution and reducing overfitting, ultimately leading to improved performance under unseen conditions.
4.3.3 Model Training
To formulate the slump estimation task as a classification problem, we discretized the continuous slump values into five categorical intervals (e.g., 150–180mm, 180–210mm), where the width of each interval reflects the ±$\pm 30$ mm tolerance specified in the KS F 4009 standard [18], along with an additional margin for human error. For intervals at the extremes (i.e., below 150 mm and above 240 mm), which are rarely encountered in practice, we did not apply explicit error margins.
We employed the ResNet 3D architecture as our backbone video classification model and trained it on labeled video clips to predict the corresponding slump category. To enhance generalization, we applied the MixUp technique, which linearly interpolates both input videos and their corresponding labels to generate augmented samples with soft targets. Under this setting, we replaced the standard cross-entropy loss with a soft-target formulation that incorporates label smoothing:
@@ 5.1 Dataset
Over a period of eight months, we constructed a dataset by replicating real-world construction site conditions. The recorded videos were segmented into 10-second intervals, resulting in $4,504$ training samples, $998$ validation samples, and $941$ test samples, as shown in Figure 4.
@@ 5.2.2 Results
Our chute detection model achieved strong quantitative results, with an mAP50-95 of 0.9945 and a precision of 0.995. The qualitative results are shown in Figure 5, where all bounding boxes exhibit confidence scores above $0.95$, indicating highly reliable detections. This demonstrates the strong and effective performance of our chute detection model. Furthermore, real-time inference is feasible, as the model achieves an average speed of $9.0$ms per frame, making it suitable for real-time applications.
@@ 5.3 Detecting Concrete Placement Location
To evaluate which chute of the concrete mixer the concrete was poured from, human annotators manually labeled the ground truth for each instance. Identifying the chute from which the concrete is poured also allows for accurate determination of the pouring start time, thereby enabling more precise quantitative evaluation. To facilitate this evaluation, we conducted experiments under three scenarios: pouring from the left-side chute, pouring from the right-side chute, and no pouring. This result is shown in Table 5.3.
As shown in the Table 5.3, the proposed system demonstrated an overall accuracy exceeding $90\%$ on average, and achieved $100\%$ accuracy in cases where concrete placement did not occur. While the chute located on the left side exhibited $100\%$ accuracy in most experimental cases, several errors were observed for the chute on the right side. This was primarily attributed to the shadow cast by the hopper cover located above the right-side chute. In particular, it was observed that after approximately 10 seconds from the start of the video, the direction of the optical flow vectors $(u,v)$ tended to follow the movement of the shadow rather than the actual flow of the concrete. However, such errors are not expected to pose significant issues in practical field applications when determining whether concrete is being placed. Additionally, in samples with slump values exceeding 240 mm, a temporary decline in accuracy was also observed for the left-side chute. This phenomenon occurred when the slump was excessively high, resulting in smooth and fluid concrete flow, which made pixel-level motion detection using optical flow more difficult. This issue was observed only in the sample with a slump value of 280 mm.

### v2

@@ Abstract
Concrete workability is essential for construction quality, with the slump test being the most common on-site method for its assessment. However, traditional slump testing is manual, time-consuming, and prone to inconsistency, limiting its applicability for real-time monitoring. To address these challenges, we propose SlumpGuard, an AI-powered, video-based system that automatically analyzes concrete flow from the truck chute to assess workability in real time. Our system enables full-batch inspection without manual intervention, improving both the accuracy and efficiency of quality control. We present the system design, the construction of a dedicated dataset, and empirical results from real-world deployment, demonstrating the effectiveness of SlumpGuard as a practical solution for modern concrete quality assurance.
@@
journal: Automation in Construction
@@
[label1]organization=Yonsei University,addressline=50 Yonsei-ro, Seodaemun-gu, city=Seoul, postcode=03772, country=Republic of Korea
@@ 1 Introduction
[…]
Concrete stands as the backbone of modern construction infrastructure, with over 14 billion cubic meters produced annually worldwide, making it the most widely consumed material after water [1]. Among the critical properties governing concrete performance, workability—primarily assessed through slump testing—represents a fundamental parameter that directly influences mixing efficiency, transportation logistics, placement operations, and compaction effectiveness on construction sites [2, 3]. However, the complex and dynamic nature of construction environments presents unprecedented challenges for maintaining consistent concrete quality control. In particular, conventional manual slump testing, still the industry standard, often fails to meet the demanding requirements of large-scale, time-sensitive projects.
The construction industry therefore continues to struggle with real-time quality assurance. Traditional slump testing remains labor-intensive, operator-dependent, and unable to provide continuous monitoring during concrete placement. Manual testing introduces delays, increases costs, and is sensitive to environmental factors, while reliance on less experienced personnel further undermines consistency [4]. Although automated approaches such as video-based slump prediction, stereo vision, depth sensing, and IoT monitoring have been explored, most are validated only in controlled or small-scale settings and face challenges such as variable lighting, dust, and equipment interference in dynamic construction environments [5, 3, 6]. Moreover, these systems often depend on specialized hardware or require significant infrastructure modifications, limiting their practicality. Consequently, there is an urgent industry need for field-deployable, cost-effective solutions capable of providing robust, real-time monitoring of concrete workability without disrupting construction workflows, addressing environmental variability, and integrating seamlessly with existing operations [7, 8].
To address the limitations of conventional slump testing in real-site applications, we present SlumpGuard, a comprehensive AI-powered system specifically engineered for real-world deployment in construction environments. Unlike manual methods that are labor-intensive, inconsistent, and unsuitable for continuous monitoring, SlumpGuard operates directly on the flow of ready-mix concrete from mixer truck chutes to enable automated, full-batch inspection of every delivery without manual intervention. Our approach introduces several key innovations: (1) a robust three-stage pipeline combining oriented object detection, optical flow analysis, and video classification, (2) advanced data augmentation strategies specifically designed to handle the environmental variability encountered in construction sites, and (3) comprehensive validation using a dataset collected over eight months from actual construction conditions rather than laboratory simulations. This video-based system not only enhances the consistency and reliability of slump measurements but also streamlines quality assurance procedures across the entire construction workflow, ultimately contributing to safer, more efficient, and higher-quality construction outcomes.
The remainder of this paper is organized as follows. Section 2 reviews existing studies on concrete workability assessment and automated monitoring approaches. Section 3 describes the dataset collection process and annotation methodology. Section 4 details the architecture of the proposed SlumpGuard system, including its three-stage pipeline and data augmentation strategies. Section 5 presents experimental results and performance evaluation under real-site conditions. Finally, Section 6 concludes the paper and outlines future research directions.
2 Background and Related Studies
@@ 2.1 Optical Flow
Optical flow is a computer vision technique used to estimate the motion of objects or scenes across video frames by analyzing the movement of pixels. It represents motion as a vector field, where each vector indicates the displacement of a pixel between consecutive frames. There are two main approaches: dense optical flow, which computes motion for every pixel, and sparse optical flow, which estimates motion only at selected feature points. While dense optical flow [9] offers a detailed view of the entire scene’s motion, it is computationally intensive. Sparse optical flow is more efficient but may overlook fine-grained motion.
@@
In this work, we adopt the Lucas-Kanade method [10], a widely used algorithm for sparse optical flow, suitable for real-time applications such as tracking concrete flow. It assumes that the intensity of a point remains constant over time:
$I(x,y,t)=I(x+\Delta x,y+\Delta y,t+\Delta t),$ (1)
and linearizes this relation using a Taylor series expansion, yielding:
$\frac{\partial I}{\partial x}u+\frac{\partial I}{\partial y}v+\frac{\partial I}{\partial t}=0,$ (2)
where $u$ and $v$ are the horizontal and vertical components of the flow. By solving this equation over a small window around each feature point, the Lucas-Kanade method efficiently estimates the local motion.
However, classical optical flow methods, including Lucas-Kanade, often struggle under challenging conditions such as illumination changes, occlusions, or non-rigid motion. To overcome these limitations, deep learning-based methods have emerged as the new state-of-the-art.
FlowNet [11] pioneered the use of convolutional neural networks for end-to-end optical flow prediction, learning motion patterns directly from synthetic datasets. Building on this idea, FlowNet2 [12], PWC-Net [13], and LiteFlowNet [14] introduced architectural improvements such as pyramid-based warping, cost volume reasoning, and hierarchical refinement. More recently, RAFT [15] achieved significant performance gains by leveraging dense, all-pairs correlations and iterative updates to produce highly accurate flow estimates.
Previous studies have applied optical flow combined with imaging techniques for slump prediction under controlled experimental settings [4, 5]. In contrast, our approach is specifically designed for robust, real-world deployment on construction sites, enabling continuous, automated slump quality monitoring under diverse and dynamic field conditions.
2.2 Object Detection
Object detection is a fundamental computer vision task widely applied in construction to automatically identify and localize key entities such as workers, equipment, and materials in images or videos. Recent studies have demonstrated its effectiveness in construction site monitoring [16, 17].
State-of-the-art detection frameworks such as Faster R-CNN [18], Mask R-CNN [19], SSD [20], and YOLO [21] offer a good balance between speed and accuracy, making them well-suited for real-time applications on construction sites.
However, in construction environments where objects like materials and equipment often appear at arbitrary angles, standard axis-aligned bounding boxes may be insufficient. In such cases, oriented object detection becomes essential. Early approaches like R2CNN [22] introduced rotation-invariant proposals, while more recent models such as Oriented R-CNN [23] and ReDet [24] employ rotation-aware architectures to improve detection accuracy. These methods are particularly effective for automated inspection tasks involving rotated or non-axis-aligned objects.
2.3 Video Classification
Video classification is a fundamental task in computer vision that involves assigning semantic labels to video clips by analyzing both spatial and temporal information inherent in the data. Recent advances in this domain have been largely driven by two primary classes of deep learning architectures: 3D convolutional neural networks (3D CNNs) and vision transformers [25] tailored for video.
3D CNNs like ResNet3D [26] extend 2D convolutions to the temporal dimension, effectively capturing spatiotemporal features and achieving strong results on benchmarks such as Kinetics-400 [27]. Transformer-based models, notably TimeSformer [28], use divided spatial and temporal self-attention to model long-range dependencies efficiently, surpassing many convolutional approaches. Further advancements include VideoSwin [29] and MViT [30], which incorporate hierarchical and multiscale attention mechanisms to improve video representation learning and classification performance.
These video classification models have also been increasingly applied in the construction industry, playing a pivotal role in streamlining quality supervision, activity recognition, and concrete workability assessment. Previous studies have also explored video-based approaches for material quality monitoring and evaluation [31, 3, 8], alongside behavior classification systems aimed at improving worker safety and productivity [32, 33].
@@ 3.2 Preprocessing and Annotation
Since we employed a stereo camera setup, the collected videos consisted of two separate streams. Therefore, as part of preprocessing, the original stereo data was divided into two individual video sequences, effectively doubling the amount of data for analysis. Considering that concrete discharges directly from the ready-mixed concrete chute, our preprocessing specifically targeted the chute region. Bounding box annotations were performed by categorizing into two classes, as shown in Figure 2, an “Unrotated Chute” representing the entire chute region with axis-aligned boxes, and a “Chute” represented by rotated bounding boxes.
@@ 4 Automation Strategy
    Assessing the discharged concrete slump to determine whether it falls within a specified range.
@@ 4.1 Chute Detection
To automatically identify when concrete is falling from the chute, we first detect the chute region in the video using an object detection model. We employ YOLOv8 [34], a state-of-the-art object detection model, known for its fast inference performance and high accuracy, making it suitable for real-time object detection. We initialize the model with weights pretrained on the DOTA dataset [35], which is designed for detecting oriented objects in aerial imagery and is well-suited for our task of detecting rotated chutes. Using this model, we simultaneously predict both axis-aligned and rotated-bounding boxes for the chutes, enabling us to detect the chute area in pouring videos. The detected chute is represented by a bounding box with parameters $(x,y,w,h,\theta)$, indicating the center coordinates, width, height, and rotation angle of the box.
@@ 4.2.1 Definition of the Bottom Edge Direction
[…]
4.2.2 Determining the Location and Timing
To make the temporal change explicit, we define the center point of the tracked chute in each frame as $(x_{t},y_{t})$ at time $t$, and as $(x_{t-1},y_{t-1})$ at the previous frame t1t-1. The optical flow vector $(u,v)$ represents the displacement from frame t1t-1 to $t$, so that $(x_{t},y_{t})=(x_{t-1}+u,y_{t-1}+v)$. The endpoints of the bottom edge of the rotated bounding box, $\mathbf{p_{1}}=(x_{1},y_{1})$ and $\mathbf{p_{2}}=(x_{2},y_{2})$, are computed.
To determine whether the tracked center point crosses the bottom edge between frames t1t-1 and $t$, we explicitly check the positions at both time steps relative to the edge. First, the line equation for the bottom edge is:
@@ 4.3.1 Data Preprocessing
We crop the unrotated-bounding box corresponding to $V$ from the full image to isolate the chute area. From the frame at time $t$, we extract a sequence of $N$ consecutive frames from the cropped region to construct a temporally continuous input.
@@ 4.3.2 Data Augmentation
Together, these augmentations help the model handle diverse visual appearances–such as changes in brightness, shadows, occlusion, and orientation–that are commonly encountered in construction sites environment. They also promote better regularization by expanding the training distribution and reducing overfitting, ultimately leading to improved performance under unseen conditions.
4.3.3 Model
We employed the ResNet 3D architecture [26] as our video classification backbone, designed to jointly capture spatial and temporal patterns. We experimented with two variants: Mixed Convolution and (2+1)D convolution. The Mixed Convolution model uses 3D convolutions in early layers for motion modeling and 2D convolutions in later layers for efficient spatial reasoning, achieving performance comparable to full 3D models with fewer parameters. The (2+1)D model factorizes 3D convolutions into separate spatial and temporal steps, enhancing optimization and model capacity.
4.3.4 Training
To formulate the slump estimation task as a classification problem, we categorized the continuous slump values into five categorical intervals (e.g. 150–180mm, 180–210mm), where the width of each interval reflects the $\pm 30$ mm tolerance specified in the KS F 4009 standard [38], along with an additional margin for human error. For intervals at the extremes (i.e. below 150 mm and above 240 mm), which are rarely encountered in practice, we did not apply explicit error margins.
To enhance generalization, we applied the MixUp technique, which linearly interpolates both input videos and their corresponding labels to generate augmented samples with soft targets. Under this setting, we replaced the standard cross-entropy loss with a soft-target formulation that incorporates label smoothing:
@@ 5.1 Dataset
Over an eight-month period, a dataset was constructed by replicating real-world construction site conditions in a controlled experimental settings. The recorded videos were segmented into 10-second intervals, resulting in $4,504$ training samples, $998$ validation samples, and $941$ test samples, as shown in Figure 4.
@@ 5.2.2 Results
Our chute detection model achieved strong quantitative results, with an mAP50-95 of 0.9945 and a precision of 0.995. The qualitative results are shown in Figure 5, where all bounding boxes exhibit confidence scores above $0.95$, indicating highly reliable detections. This demonstrates the strong and effective performance of our chute detection model. Furthermore, real-time inference is feasible, as the model achieves an average speed of 9.09.0ms per frame, making it suitable for real-time applications.
@@ 5.3 Detecting Concrete Placement Location
To evaluate which chute of the mixer truck the concrete was poured from, human annotators manually labeled the ground truth for each instance. Identifying the chute from which the concrete is poured also allows for accurate determination of the pouring start time, thereby enabling more precise quantitative evaluation. To facilitate this evaluation, we conducted experiments under three scenarios: pouring from the left-side chute, pouring from the right-side chute, and no pouring. This result is shown in Table 1.
As shown in the Table 1, the proposed system demonstrated an overall accuracy exceeding $90\%$ on average, and achieved $100\%$ accuracy in cases where concrete placement did not occur. While the chute located on the left side exhibited $100\%$ accuracy in most experimental cases, several errors were observed for the chute on the right side. This was primarily attributed to the shadow cast by the hopper cover located above the right-side chute. In particular, it was observed that after approximately 10 seconds from the start of the video, the direction of the optical flow vectors $(u,v)$ tended to follow the movement of the shadow rather than the actual flow of the concrete. However, such errors are not expected to pose significant issues in practical field applications when determining whether concrete is being placed. Additionally, in samples with slump values exceeding 240 mm, a temporary decline in accuracy was also observed for the left-side chute. This phenomenon occurred when the slump was excessively high, resulting in smooth and fluid concrete flow, which made pixel-level motion detection using optical flow more difficult. This issue was observed only in the sample with a slump value of 280 mm.